40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Wed Aug 2 03:35:18 2017, Zach, Electronics, SiFi, Current driver range increased to 200 mA SiFi_current_driver2_coarse_response_8_1_17.pngSiFi_current_driver2_fine_response_8_1_17.pngR31.jpg
    Reply  Wed Aug 2 08:31:58 2017, Johannes, Electronics, SiFi, Current driver range increased to 200 mA 
       Reply  Wed Aug 2 17:23:48 2017, Zach, Electronics, SiFi, Current driver range increased to 200 mA SN104978_PvsI.pdf
    Reply  Thu Aug 3 18:06:05 2017, Zach, Update, SiFi, More system characterization measurements (laser drivers, heads) SiFi_laser_PvsI_8_2_17.png
Message ID: 1658     Entry time: Thu Aug 3 18:06:05 2017     In reply to: 1652
Author: Zach 
Type: Update 
Category: SiFi 
Subject: More system characterization measurements (laser drivers, heads) 

DC current reponse of laser driver #1

To help keep things moving along*, Johannes gave me a spare laser driver box that already has a 200-mA range (S1600246 - the new SiFi "Driver #1").

Here are the knob -> current responses of this board:

 

* I am still planning to get the other unit (S1500278) back up and running when I have time.

 

Laser head P vs. I curves

Next I measured the output power of the laser heads as a function of drive current. Here they are compared with the manufacturer's data (inferred by eye from crappy plots):

Clearly, the outputs begin to underperform the specs at higher drive, but not by a staggering amount. These units are a couple years old, so I don't believe(?) this reduction in output is unexpected.

 

Quote:

One of the things I would like to have is a detailed characterization of my setup, both for easy reference and for my thesis. A natural place to start is the laser drivers (D1200719).

[...]

Here are the knob -> current responses of current driver #2 (S1500267) after the modification:

[...]

So... where's #1? Alas, I actually did this mod on #1 on Monday night, but the thing didn't behave properly after it and I'm still looking into it...

 

 

ELOG V3.1.3-